OpenStudy (anonymous):

A rectangular box is x feet long and x feet wide. The volume of the box is (4x8 + 3x6) cubic feet. What polynomial represents the height of the box in feet? A. 4x + 3 B. 4x6 + 3x4 C. 4x8 + 3x6 D. 4x7 + 3x5

OpenStudy (anonymous):

@turtleluver @SolomonZelman

OpenStudy (anonymous):

Height = V/lw

jigglypuff314 (jigglypuff314):

volume of box: (length)(width)(height) so 4x^8 + 3x^6 = (x)(x)(height) so to find height factor out (x)(x) from 4x^8 + 3x^6

OpenStudy (anonymous):

Height = (4x^8 + 3x^6)/x^2

OpenStudy (anonymous):

what is that equal to?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

right?

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

yay! :D ok

OpenStudy (anonymous):